OpenStudy (anonymous):

Find the equation of a line that contains the point (2,-3) with a slope of 1/6.

OpenStudy (the_fizicx99):

y-y1=m(x-x1)

OpenStudy (anonymous):

put [m=1/6 , y=-3 , x=2] in \[m=\frac{ y-y1 }{ x-x1}\]

OpenStudy (anonymous):

@Abdulhameed

OpenStudy (the_fizicx99):

y-(-3) = 1/6(x-2) y + 3 = 1/6(x -2) y + 3 = 1/6x - 1/3 +3 +3 Solve: -1/3 + 3/1 Then put it in y = mx + b You only need the y intercept